OpenStudy (anonymous):

What is the y-intercept of the line 3x + 4y = 16?

jimthompson5910 (jim_thompson5910):

Plug in x = 0 and solve for y 3x + 4y = 16 3(0) + 4y = 16 0 + 4y = 16 4y = 16 I'll let you finish up

OpenStudy (anonymous):

thanks your a life saver!(:

jimthompson5910 (jim_thompson5910):

you're welcome